OpenStudy (anonymous):

Solve for y in the following system of equations. 2x - 2y = 8 -2x + 8y = 4

OpenStudy (anonymous):

ThANK YOU SO MUCH

OpenStudy (anonymous):

Same thing: 2x + (-2y) = 8 + -2x + 8y = 4 ------------------ 0x + 6y = 12 6y = 12 y = 2

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2x=8+2y x = 4+y -2(4+y)+8y=4 -8-2y+8y=4 6y=12 y=2 x=4+y x=6